OpenStudy (anonymous):

What are the solutions of the quadratic equation? 4x^2-9x-9=0 And can you show me the steps?

OpenStudy (anonymous):

\[4x ^{2} -9x -9 = 4x ^{2} -12x + 3x -9=4x(x-3)+3(x-3) =(4x+3)(x+3)\]

OpenStudy (anonymous):

i splited it as (-12) and 3 because 4*(-9)= -36 which can also be written as -12*3=-36

OpenStudy (anonymous):

and since term with "x" is negative .. therefore 12 is taken as negative.... if it would have been positive.... 3 would have been taken negative...

OpenStudy (zpupster):

(4x + 3)(x - 3) is his final he made a mistake in writing the final answer x-3 instead of x+3